### v1

Abstract
The value of second-order methods lies in the use of curvature information. Yet, this information is costly to extract and once obtained, valuable negative curvature information is often discarded so that the method is globally convergent. This limits the effectiveness of second-order methods in modern machine learning. In this paper, we show that second-order and second-order-like methods are promising optimizers for neural networks provided that we add one ingredient: negative step sizes. We show that under very general conditions, methods that produce ascent directions are globally convergent when combined with a Wolfe line search that allows both positive and negative step sizes. We experimentally demonstrate that using negative step sizes is often more effective than common Hessian modification methods.
1 Introduction
Modern machine learning typically involves training deep neural networks and is thus a high-dimensional and non-convex problem. Non-convexity makes learning difficult for second-order methods that are attracted to saddle points and local maxima. By and large, gradient based methods such as gradient descent with momentum, Adam (Kingma and Ba, 2014) and RMSprop (Tieleman and Hinton, 2012) are seen as the only viable optimizers in machine learning. This is a source of frustration. Newton’s method and second-order-like methods, such as limited-memory quasi-Newton (QN) methods, excel in traditional machine learning but can fail to even converge in the deep learning setting. First-order methods converge but slowly, as good convergence progress can only be expected if curvature information of the loss landscape is used. We want optimizers that use negative curvature information but these methods may yield search directions that point the wrong way.
Many variants of second-order methods exist that ensure global convergence by enforcing a positive-definite second-order approximation. This approach may not be accurate for non-convex problems where the Hessian may have negative eigenvalues. Other methods, such as the dogleg method (Powell, 1970) and trust-region methods, combine gradient descent (GD) with a second-order method to ensure progress. Yet, for problems that are ill-conditioned and non-convex, GD ensures progress but at an extremely slow rate (Martens, 2010) and the second-order direction could point uphill or to saddle points. In such cases, simply taking the negative of the second-order direction may be a better choice, as this points downhill while maintaining all second-order information (see Figure 1).
To the best of our knowledge, using negative step sizes is relatively unexplored in optimization. We found one mention of negative step sizes in an analysis of the Fletcher-Reeves Fletcher and Reeves (1964) non-linear conjugate gradient method Dai (1999) for differentiable and Lipschitz-smooth objective functions.
1.1 Contribution
This paper examines the largely unexplored role of negative step sizes in optimization and show that taking a backward step is a computationally inexpensive way to incorporate negative curvature information. Our experiments suggest that a quasi-Newton (QN) method, symmetric rank 1 (SR1) Conn et al. (1991), combined with negative step sizes, may be an overlooked method useful for training neural networks. This is achieved without using common approaches to “fix” second-order methods [^1] In this paper, we refer to second-order-like methods that approximate the Hessian, such as QN methods, as “second-order methods”. This is technically incorrect because these methods do not compute second derivatives. This misuse of terminology is for readability. such as Hessian modifications. We also show that extending the Wolfe line search to both positive and negative step sizes ensures that second-order methods satisfy the Zoutendijk condition. Thus, second-order methods combined with good step size choices are globally convergent even in non-convex settings.
Figure 1: In this two-dimensional non-convex minimization problem, the negative of the Newton’s direction (green) is a better search direction than the gradient descent direction (blue), Newton’s direction (red) or any convex combination of both.
2 Common approaches for globally convergent second-order methods
Two major drawbacks of using second-order methods to train neural networks are (a) the computation and storage cost of the (approximate) Hessian and, (b) non-convergence. To some extent, the first issue is addressed by QN methods and their limited memory versions Nocedal and Wright (2006). In this section, we summarize some common approaches to the second issue of non-convergence.
Hessian modifications
There are many variants of second-order method that modify the (approximate) Hessian for positive definiteness and thus ensure a descent direction (see Sections 3.4 of Nocedal and Wright (2006)). A spectral decomposition will find the negative eigenvalues. One could then zero out the negative eigenvalues, flip the sign of the negative eigenvalues Paternain et al. (2019) or shift the Hessian by subtracting a diagonal matrix that contains the most negative eigenvalue of the Hessian (also known as damping). These modifications, however, require an expensive eigenvalue decomposition and may no longer reflect accurate curvature information. Removing negative eigenvalues could also be detrimental for high-dimensional non-convex problems because following paths of negative curvature is a way to escape saddle points Choromanska et al. (2015); Moré and Sorensen (1979).
Figure 2: Using negative step sizes in training neural networks. The plots show training error by iteration for neural networks with 1-, 2- and 3- hidden layers on the heart dataset. We compare full QN methods BFGS (yellow) and SR1 (blue and green) with GD (red) and Adam (black). SR1 is non-convergent with positive-only step sizes (blue) but outperforms when step sizes are allowed to be negative (green). The outperformance appears more pronounced in deeper networks.
Trust-region approaches
Trust region (TR) methods is an option for obtaining global convergence and, unlike line searches, they work with methods that produce ascent directions Conn et al. (2000). Newton’s and QN methods often employ a TR approach that minimizes the model function over a two-dimensional subspace Sorensen (1982); Byrd et al. (1988); Schultz et al. (1985); Ramanmurthy and Duffy (2016). Saddle-free Newton’s method Pascanu et al. (2014); Dauphin et al. (2014), Newton’s method with generalized TR, was proposed as a version of Newton’s method that can escape saddle points. However, TR is comparable to damping Moré and Sorensen (1983) and thus loses curvature information in the same way as Hessian modifications.
Cubic regularization
Cubic-regularization (CR) uses an additional third-order term so that the second-order method is not attracted to local maxima and saddle points. Originally introduced for Newton’s method Nesterov and Polyak (2006), CR was recently extended to QN methods (Kamzolov et al., 2023; Benson and Shanno, 2018). However, CR generally requires special solvers for its third-order subproblem.
Positive definite-enforcing updates
Some QN methods, such as BFGS (Broyden, 1967; Fletcher and Reeves, 1964; Goldfarb, 1970; Shanno, 1970), are designed to guarantee positive definiteness of the Hessian approximation at every update step as long as the step size used satisfies the Wolfe conditions (Wolfe, 1969). It has been suggested, however, that QN variants that do not enforce positive definiteness, such as SR1, converge faster by adhering closer to the true Hessian (Conn et al., 1991; Khalfan et al., 1993). Other updates that enforce positive definiteness include natural gradient methods Amari (1998) that use the covariance matrix of the gradients instead of the Hessian, Gauss-Newton methods and Kronecker-factored approximate curvature (KFAC) Martens and Grosse (2015). These methods, however, may also sacrifice accurate curvature information by enforcing positive definiteness.
3 Global convergence of ascent-descent methods
Even though negative step sizes may result in a larger decrease on some iterations, we may be concerned that the algorithm might not converge. In this section, we give a general result showing that methods that may produce ascent directions and that allow negative step sizes converge under very general conditions. Our goal is to minimize a function $f:{\mathbb{R}}^{n}\rightarrow{\mathbb{R}}$ without constraints. $f$ is assumed to be twice differentiable, Lipschitz-smooth and bounded below. We use a deterministic iterative algorithm that, on the $k$th iteration, calculates a search direction ${\bm{p}}_{k}$ and step size $\alpha_{k}$, and updates its iterate ${\bm{x}}$ by
${\bm{x}}_{k+1}={\bm{x}}_{k}+\alpha_{k}{\bm{p}}_{k}.$ (1)
Dataset size GD Adam l-BFGS l-SR1 l-SR1+damp l-SR1+Wolfe±$\pm$
a1a $(1605\times 119)$ 0.217 0.203 0.315 0.500 0.214 0.199
a9a $(32561\times 123)$ 0.220 0.212 0.332 0.330 0.216 0.211
colon-cancer $(62\times 2000)$ 0.006 0.046 0.500 0.183 DNF 0.003
gisette $(6000\times 5000)$ 0.139 0.171 0.395 0.395 DNF 0.050
heart $(270\times 13)$ 0.455 0.382 0.388 0.237 0.239 0.237
ijcnn1 $(35000\times 22)$ 0.132 0.120 0.131 0.165 0.130 0.120
ionosphere $(351\times 34)$ 0.201 0.118 0.201 0.500 0.199 0.170
leukemia $(38\times 7129)$ 0.028 0.001 0.160 0.160 DNF 0.064
madelon $(2000\times 500)$ 0.500 0.500 0.490 0.396 0.382 0.396
mushrooms $(8124\times 112)$ 0.040 0.018 0.166 0.500 0.019 0.009
splice $(1000\times 60)$ 0.270 0.260 0.485 0.500 0.262 0.195
svmguide3 $(1243\times 22)$ 0.295 0.276 0.374 0.374 0.267 0.260
w1a $(2477\times 300)$ 0.088 0.059 0.200 0.200 0.071 0.063
w8a $(49749\times 300)$ 0.096 0.066 0.193 0.193 0.079 0.071
Table 1: Comparing the use of negative step sizes with damping, a common Hessian modification, for limited-memory QN methods. The experiment was run across several datasets fitted with a neural network. The lowest training error achieved in each dataset is highlighted in bold. Limited-memory SR1 with positive only step sizes (l-SR1) often does not converge. Damping (l-SR1+damp) helps with convergence but is an expensive operation and in many cases did not finish training in the allocated time (shown as DNF). Using negative step sizes (l-SR1+Wolfe±$\pm$) showed good performance even against a state-of-the-art optimizer such as Adam.
Given our iterate ${\bm{x}}_{k}$ and a search direction ${\bm{p}}_{k}$, we can rewrite $f$ as a function of the step size $\phi(\alpha)=f({\bm{x}}_{k}+\alpha{\bm{p}}_{k})$. The directional derivative is defined as its derivative with respect to $\alpha$, i.e. $\phi^{\prime}(\alpha)=\nabla f({\bm{x}}_{k}+\alpha{\bm{p}}_{k})^{\top}{\bm{p}}_{k}$. We refer to ${\bm{p}}_{k}$ as an ascent or a descent direction when $\phi^{\prime}(0)$ is positive or negative respectively. We say that a method is globally convergent if it finds a stationary point using update rule 1 starting from any initial estimate ${\bm{x}}_{0}$. Further details of notation, definitions and assumptions are given in Appendix A.
It is known that methods that produce descent-only directions with step sizes satisfying the Wolfe conditions Wolfe (1969) are globally convergent for a loss function $f$ that satisfies the assumptions above Nocedal and Wright (2006). The proposition below extends this result to methods that produce both ascent and descent directions, and that satisfy a variant of the Wolfe conditions that considers both positive and negative step sizes. Global convergence holds with the looser requirement that ${\bm{p}}_{k}$ is not orthogonal to the gradient direction.
Proposition 3.1.
Consider any algorithm with iteration of the form equation 1 where ${\bm{p}}_{k}$ is direction that is not orthogonal to the gradient direction, and non-zero step sizes $|\alpha_{k}|\geq\epsilon>0$ satisfy the Wolfe conditions (6 and 7). Suppose that $f$ is bounded below in $\mathcal{R}^{n}$ and that $f$ is continuously differentiable in an open set $\mathcal{N}$ containing the level set $\triangleq$$\mathcal{L}\triangleq\{{\bm{x}}:f({\bm{x}})\leq f({\bm{x}}_{0})\}$, where ${\bm{x}}_{0}$ is the initial iterate. Assume also that the gradient $\nabla f$ is Lipschitz continuous on $\mathcal{N}$. Then, the algorithm satisfies Zoutendijk’s condition (4) is therefore globally convergent (2).
The proof of Proposition 3.1 is given in Appendix B. A modified Wolfe line search that also considers negative step sizes could be easily implemented. We give one possible implementation, which we call Wolfe±$\pm$, in Appendix C.
4 Experiments
4.1 QN methods and neural network training
In this experiment, we looked at full QN methods and their performance on neural network training. All methods use a line search that gives step sizes satisfying the Wolfe conditions (p). SR1 is further combined with Wolfe±$\pm$ (p+n) that uses negative step sizes when ${\bm{p}}_{k}$ is not a descent direction. QN methods are initialized with the identity matrix. Methods are run for a maximum of 50 iterations. Figure 2 shows training error by iteration.
Because the objective is non-convex, the true Hessian at iterate ${\bm{x}}_{k}$ may not be positive definite. Thus, SR1 produces Hessian approximations that are not positive definite and a ${\bm{p}}_{k}$ that does not point downhill. With a regular line search, SR1 diverges within the first few steps (blue). With negative step sizes, SR1 becomes the an effective optimizer. This effect appears to be more pronounced as the number of network layers increase. BFGS combined with a Wolfe line search guarantees positive definiteness in its Hessian update and thus never diverges even with positive only step sizes. Yet, a positive definite Hessian may not accurately capture the optimization landscape. The lack of negative curvature information in the BFGS Hessian approximation may be the reason its training error is higher than that of SR1 after the maximum number of iterations is reached.
For further details, please refer to Appendix D.
4.2 Limited-memory QN methods and neural network training
Table 1 shows the result of using two limited-memory QN methods, l-BFGS and l-SR1, to train a neural network across different datasets. Gradient descent (GD) and Adam are plotted for comparison. The purpose of this experiment is to compare two globalization strategies: damping and negative step sizes. Damping requires an eigenvalue decomposition and is therefore an expensive operation. Thus, in many cases where the dataset has a large number of features and the Hessian is large (e.g. colon-cancer), l-SR1 combined with damping did not finish in the allocated amount of time (shown as DNF). l-BFGS combined with a Wolfe line search does not require damping because its Hessian approximations are guaranteed to be positive definite. Yet, a lack of negative curvature information may hurt its effectiveness in training neural networks, and it tends to perform worse than l-SR1 with negative step sizes.
Acknowledgements
Betty Shea is funded by an NSERC Canada Graduate Scholarship. The work was partially supported by the Canada CIFAR AI Chair Program and NSERC Discovery Grant RGPIN-2022-036669.
Appendix A Notation and further background details
We say an algorithm is globally convergent if it produces a sequence of gradients that converge to zero, or
$\liminf_{k\rightarrow\infty}\|\nabla f({\bm{x}}_{k})\|=0.$ (2)
The angle $\theta_{k}$ between search direction ${\bm{p}}_{k}$ and the steepest descent direction $-\nabla f({\bm{x}}_{k})$ is given by
$\cos\theta_{k}=\frac{-\nabla f({\bm{x}}_{k})^{\top}{\bm{p}}_{k}}{\|\nabla f({\bm{x}}_{k})\|\|{\bm{p}}_{k}\|}.$ (3)
Equation (3) is sometimes referred to as the cosine similarity between the ${\bm{p}}_{k}$ and the steepest descent direction. Search direction ${\bm{p}}_{k}$ is not orthogonal to the gradient if $|\cos\theta_{k}|\geq\delta>0$.
Zoutendijk’s condition is satisfied if
$\sum_{k\geq 0}\cos^{2}\theta_{k}\|\nabla f({\bm{x}}_{k})\|^{2}<\infty,$ (4)
which implies that
$\cos^{2}\theta_{k}\|\nabla f({\bm{x}}_{k})\|^{2}\rightarrow 0.$
If $f$ has a Lipschitz continuous gradient on open set $\mathcal{N}$, then there exists a constant $L>0$ such that
$\|\nabla f({\bm{x}})-\nabla f(\tilde{{\bm{x}}})\|\leq L\|{\bm{x}}-\tilde{{\bm{x}}}\|\text{, for all }{\bm{x}},\tilde{{\bm{x}}}\in\mathcal{N}$ (5)
A line search is an auxiliary method that finds a step size $\alpha_{k}$ along ${\bm{p}}_{k}$ that has desirable properties. For example, for constants $c_{1}\in(0,1)$ and $c_{2}\in(c_{1},1)$ and descent direction ${\bm{p}}_{k}$, a step size $\alpha_{k}>0$ that satisfies the Wolfe condition has two desirable properties. Firstly, $\alpha_{k}$ guarantees making progress in minimizing $f$, or
$f({\bm{x}}_{k}+\alpha_{k}{\bm{p}}_{k})\leq f({\bm{x}}_{k})+c_{1}\alpha_{k}\nabla f({\bm{x}}_{k})^{\top}{\bm{p}}_{k}.$ (6)
Equation 6 is also known as the Armijo condition. Secondly, $\alpha_{k}$ does not lie too far from an optimal step size choice.
$0\geq\nabla f({\bm{x}}_{k}+\alpha_{k}{\bm{p}}_{k})^{\top}{\bm{p}}_{k}\geq c_{2}\nabla f({\bm{x}}_{k})^{\top}{\bm{p}}_{k}.$ (7)
Equation equation 7 is also known as the curvature condition. Taken together, equations equation 6 and equation 7 are often referred to as the Wolfe conditions. Note that this standard curvature condition assumes that ${\bm{p}}_{k}$ is a descent direction and thus $\nabla f({\bm{x}}_{k})^{\top}{\bm{p}}_{k}<0$. In our case where ${\bm{p}}_{k}$ is an ascent direction, the curvature condition becomes
$0\leq\nabla f({\bm{x}}_{k}+\alpha_{k}{\bm{p}}_{k})^{\top}{\bm{p}}_{k}\leq c_{2}\nabla f({\bm{x}}_{k})^{\top}{\bm{p}}_{k}.$ (8)
Appendix B Proof of Proposition 1
Proof B.1.
The proof of Theorem 3.1 follows closely that of Theorem 3.2 in Nocedal and Wright (2006), which considers only the case of descent directions paired with positive step sizes. Here we extend the results to ascent directions ${\bm{p}}_{k}$ paired with negative step sizes $\alpha_{k}$.
Combining update rule (1) and curvature condition (8) where $\nabla f({\bm{x}}_{k})^{\top}{\bm{p}}_{k}>0$ and $\alpha_{k}<0$ gives
$\left(\nabla f({\bm{x}}_{k})-\nabla f({\bm{x}}_{k+1})\right)^{\top}{\bm{p}}_{k}\geq(1-c_{2})\nabla f({\bm{x}}_{k})^{\top}{\bm{p}}_{k}.$ (9)
Lipschitz continuous gradient gives
$\|\nabla f({\bm{x}}_{k})-\nabla f({\bm{x}}_{k+1})\|\leq L\|{\bm{x}}_{k}-{\bm{x}}_{k+1}\|=L\|-\alpha_{k}{\bm{p}}_{k}\|=|-\alpha_{k}|L\|{\bm{p}}_{k}\|=-\alpha_{k}L\|{\bm{p}}_{k}\|$
and
$\left(\nabla f({\bm{x}}_{k})-\nabla f({\bm{x}}_{k+1})\right)^{\top}{\bm{p}}_{k}\leq-\alpha_{k}L\|{\bm{p}}_{k}\|^{2}.$ (10)
Combining (9) and (10)
$(1-c_{2})\nabla f({\bm{x}}_{k})^{\top}{\bm{p}}_{k}\leq-\alpha_{k}L\|{\bm{p}}_{k}\|^{2}.$
Rearranging gives
$\alpha_{k}\leq-\frac{1-c_{2}}{L}\frac{\nabla f({\bm{x}}_{k})^{\top}{\bm{p}}_{k}}{\|{\bm{p}}_{k}\|^{2}}\leq 0$ (11)
where the middle term in inequality (11) is negative because $10<c_{1}<c_{2}<$ 1 and $\nabla f({\bm{x}}_{k})^{\top}{\bm{p}}_{k}>0$. Using (11) with the sufficient decrease condition (6) gives
$f({\bm{x}}_{k+1})$ $\leq f({\bm{x}}_{k})+c_{1}\alpha_{k}\nabla f({\bm{x}}_{k})^{\top}{\bm{p}}_{k}$
$\leq f({\bm{x}}_{k})-\frac{c_{1}(1-c_{2})}{L}\frac{\left(\nabla f({\bm{x}}_{k})^{\top}{\bm{p}}_{k}\right)^{2}}{\|{\bm{p}}_{k}\|^{2}}$
$=f({\bm{x}}_{k})-\frac{c_{1}(1-c_{2})}{L}\left(\frac{\left(\nabla f({\bm{x}}_{k})^{\top}{\bm{p}}_{k}\right)}{\|\nabla f({\bm{x}}_{k})\|\|{\bm{p}}_{k}\|}\right)^{2}\|\nabla f({\bm{x}}_{k})\|^{2}$
$=f({\bm{x}}_{k})-c\cdot\cos^{2}\theta_{k}\|\nabla f({\bm{x}}_{k})\|^{2}$
where the last step uses (3) and $c=\frac{c_{1}(1-c_{2})}{L}$. Summing across iterations 0 to $k$,
$f({\bm{x}}_{k+1})\leq f({\bm{x}}_{0})-c\sum_{j=0}^{k}\cos^{2}\theta_{j}\|\nabla f({\bm{x}}_{j})\|^{2}$
Because $f$ is bounded below, we know that $f({\bm{x}}_{0})-f(x_{k})$ is less than a positive constant for all $k$. Taking limits gives
$\sum_{k=0}^{\infty}\cos^{2}\theta_{k}\|\nabla f({\bm{x}}_{k})\|^{2}<\infty$
and thus satisfies the Zoutendijk condition, and implies that
$\cos^{\theta}_{k}\|\nabla f({\bm{x}}_{k})\|^{2}\rightarrow 0$
Because $|\cos\theta_{k}|\geq\delta>0$ for all $k$, then $\cos^{2}\theta_{k}\geq\delta^{2}>0$ for all $k$. This further implies that
$\lim_{k\rightarrow\infty}\|\nabla f({\bm{x}}_{k})\|=0$
and thus the algorithm is globally convergent.
Appendix C Generalized Wolfe conditions
Suppose we have an implementation of the standard Wolfe line search (Nocedal and Wright (2006) Section 3.5). Then Wolfe±$\pm$ can be implemented trivially as in Algorithm C. First we check the sign of the directional derivative. If ${\bm{p}}_{k}$ is a descent direction, then we continue by calling the standard Wolfe line search. If ${\bm{p}}_{k}$ is an ascent direction, then the standard Wolfe line search is called with $-{\bm{p}}_{k}$ as the direction and the step size returned is negated afterwards.
[…]
Appendix D Further details on experiments
Figure 3 plots the step sizes and cosine similarity of the search direction ${\bm{p}}_{k}$ with the steepest descent direction of the different optimizers in the experiments discussed in Section 4.1.
Figure 3: Step sizes (top row) and cosine similarity (bottom row) of different optimizers for neural networks with 1, 2 and 3 hidden layers. SR1 often produces ascent directions and this is dealt with effectively by using negative step sizes.